• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 16:01:38 on February10,2009

    Entry number 260318

    keyword=Day shift summary


    Shift workers: Ole Hansen, Dustin McNulty (TO), Xiaochao Zheng (3rd)

    Main events:

    10:30 Access to hall (for polarity changes etc.) finished. BB is NEG, LHSR is POS.
    10:50 Beam checkout. Found raster "4.2 x 6.2" at MCC gives us about 4 x 4 mm on target.
    11:25 BigBite optics study. BB sieve IN. Raster on/off. Carbon and hydrogen targets. (Xin)
    13:50 Controlled access to remove BigBite sieve slit.
    14:25 Finishing BB optics study with sieve out.
    14:35 Start BigBite chamber HV scan (Kalyan).
    15:20 BigBite momentum and preshower/shower calibration. Hydrogen target. In progress at endof shift. (Kalyan)

    Data:

    Beam checkout
    Run Time Tgt I #events comments
    1244 1056-1103 BeO/C 2 92k raster 2x2 (MCC)
    1245 1104-1111 BeO/C 2 91k raster 2.1x3.1 (MCC) -> 2x2 (spot++)

    Optics study. BB neg w/sieve, LHRS pos.
    1246 1124-1126 BeO/C 5 76k raster off
    1247 1127-1142 BeO/C 5 774k "
    1248 1146-1147 BeO/C 5 66k raster 4x4
    1249 1148-1200 BeO/C 5 705k "
    1250 1205-1219 BeO/C 5 686k raster off, BB magnet OFF
    1251 1239-1324 H2 15 2M raster 4x4, BB magnet ON
    1252 1325-1350 H2 15 530k "
    controlled access to remove BB slieve
    1253 1423-1424 BeO/C 5 100k junk, wrong BB HV, BB w/o sieve
    1254 1426-1434 BeO/C 5 501k
    change DAQ from TWOARM_FBADC to TWOARM_FB (Cherenkov readout change)

    BigBite HV/current/rate scan
    1255 1443-1451 BeO/C 6 492k
    1256 1500-1506 BeO/C 7 420k
    1257 1511-1515 BeO/C 4 280k

    BigBite momentum and preshower/shower calibration
    1258 1524-1541 H2 15 1M
    1259 1541- H2 15 in progress at end of shift

    H2 target means: reference cell filled with hydrogen @ 140psi